# Hierarchy of timescales in a disordered spin-$1/2$ XX ladder

### v2

@@ III.1.2 Numerical extraction of $t_{\mathrm{RMT}}$
A straightforward approach to extract $\tau_{\mathrm{RMT}}$ from $\overline{K}{(\tau)}$ involves comparing it with the spectral form factor of a RMT eigenvalue distribution, such as the Gaussian Orthogonal Ensemble (GOE). This comparison leads to the following criterion function [22]:
@@
$\Delta K{(\tau>0)}<\epsilon_{\Delta K}\,.$ (7)
For the comparison with other relaxation timescales, the RMT time must be expressed in real units. This can be achieved by multiplying it by the real Heisenberg time, $t_{\mathrm{H}}$, as follows:
@@
To eliminate the variations in the density of states $\rho{(E;\boldsymbol{\mathcal{W}})}$ due to the ensemble of the various disorder realizations appearing in the SFF computations, we unfold the energy spectrum of each realization $\boldsymbol{\mathcal{W}}$, which also makes comparing the numerical results with RMT easier and thus highlights the universal RMT behavior [121, 61, 122, 123]. As a filter function, we use a Gaussian distribution with a given filter width factor $\eta$. However, for some disorder realizations, even after unfolding the energy spectrum, $\tau_{\mathrm{H}}$ may shift from $2\pi$ because of the spectral edge effects [124]. To prevent this shifting, we renormalize $\tau_{\mathrm{H}}$ in each disorder realization by applying the filter function to the mean-level spacing of the unfolded spectrum. We explain this shift-fixing procedure in Appendix B.3 and further details regarding the unfolding and filtering procedures can be found in Appendices B.1 and B.2, respectively.
The choices for $\epsilon_{\Delta K}$ and $\eta$ quantitatively affect the result for $t_{\mathrm{RMT}}$. However, the $L^{2}$ scaling of $t_{\mathrm{RMT}}$ observed in our model remains robust as long as $\epsilon_{\Delta K}$ is chosen to be sufficiently large to surpass the fluctuations that occur just prior to the onset of RMT behavior in $\overline{K}_{\mathrm{RA}}{(t)}$, and $\eta$ falls within a reasonable range, typically between $0.2$ and $1.0$. As part of our analysis, we estimate the RMT time, $t_{\mathrm{RMT}}$, for the models by varying the threshold value $\epsilon_{\Delta K}$ over a range of $0.035$ to $0.2$, while keeping the parameter $\eta$ fixed at $0.5$. Generally, we search for the range of pairs $(\eta,\varepsilon_{\Delta K})$ that yield the same $L$-dependence of $t_{\mathrm{RMT}}$, from which we then extract $t_{\mathrm{RMT}}$ up to this systematic uncertainty. Thus the exact prefactor of $t_{\mathrm{RMT}}$ is only fixed within these limitations.
In our results, we obtain $\overline{K}_{\mathrm{RA}}{(\tau)}$ in Eq. (6) by time-averaging $\overline{K}{(\tau)}$ with a running average window in $\log_{10}$ scale, $\delta{(\log_{10}\tau)}$, of $0.025/J_{\parallel}$ while the discrete time step in $\log_{10}$ scale, $\Delta{(\log_{10}\tau)}$, is $0.001/J_{\parallel}$.
@@ III.2 Spin- and energy-Thouless times
Our main interest is in a direct comparison of Thouless times, extracted from transport, to the RMT time. We recall that the Thouless time $t_{\mathrm{Th}}$ is the average time for a local excitation to spread through a finite system with length $L$:
@@
here assuming diffusion. The symbol $\mu$ indicates the transport channel, such as spin, particle, charge, or energy, and $D^{(\mu)}$ is the diffusion constant for the corresponding transport channel in the thermodynamic limit. In Eq. (9), we divide $L$ by $2\pi$ to obtain the smallest wavenumber $k_{0}=2\pi/L$. This wavenumber also appears in the autocorrelation function of the local densities, $c^{(\mu)}{(k,t)}$. For late times, this autocorrelation function decays exponentially at wavenumber $k=k_{0}$ [125], with a decay rate given by:
@@ III.2.1 Transport quantities
$\begin{cases}Q^{(\mathrm{S})}=\sum_{\ell,i}S^{z}_{\ell,i}=:S^{z}\\ Q^{(\mathrm{E})}{(\boldsymbol{\mathcal{W}})}=H{(\boldsymbol{\mathcal{W}})}\end{cases}\,$, (11)
where $Q^{(\mathrm{S})}$ is the total magnetization in the $z$-direction of the spins, and $Q^{(\mathrm{E})}$ is the total energy.
@@
$\dot{q}_{\ell}{(\boldsymbol{\mathcal{W}})}$ $=$ $i\commutator{H{(\boldsymbol{\mathcal{W}})}}{q^{(\mu)}_{\ell}{(\boldsymbol{\mathcal{W}})}}\,$, (12)
$=$ $j^{(\mu)}_{\ell-1}{(\boldsymbol{\mathcal{W}})}-j^{(\mu)}_{\ell}{(\boldsymbol{\mathcal{W}})}\,$, (13)
where $q^{(\mu)}_{\ell}{(\boldsymbol{\mathcal{W}})}$ is the local density of transport quantity at the $\ell$-rung, so that $Q^{(\mu)}{(\boldsymbol{\mathcal{W}})}=\sum_{\ell}q_{\ell}{(\boldsymbol{\mathcal{W}})}$. From this equation, we can derive expressions for the local current operator $j^{(\mu)}_{\ell}{(\boldsymbol{\mathcal{W}})}$, which are essential for characterizing the transport properties of the system. As expected, these current operators may differ depending on the choice of local density.
III.2.2 Local densities of transport quantities & local current operators
@@
$j_{(\ell,\ell^{\prime}),(i,i^{\prime})}\coloneqq\frac{1}{2i}\quantity(S^{+}_{\ell,i}S^{+}_{\ell^{\prime},i^{\prime}}-S^{-}_{\ell,i}S^{-}_{\ell^{\prime},i^{\prime}})\,.$ (15)
Similarly, the local energy current $j_{\ell}^{\mathrm{(E)}}$ can be obtained by selecting the total energy as the transport quantity, i.e., $Q^{\mathrm{(E)}}{(\boldsymbol{\mathcal{W}})}=\sum_{\ell}q^{\mathrm{(E)}}_{\ell}{(\boldsymbol{\mathcal{W}})}$, where $q^{\mathrm{(E)}}_{\ell}{(\boldsymbol{\mathcal{W}})}$ is the local energy density. For our model, we express $q^{\mathrm{(E)}}_{\ell}{(\boldsymbol{\mathcal{W}})}$ as the sum of three terms:
$q^{\mathrm{(E)}}_{\ell}{(\boldsymbol{\mathcal{W}})}\coloneqq q^{\mathrm{(E)}}_{\parallel,\ell}+q^{\mathrm{(E)}}_{{}_{\perp},\ell}+q^{\mathrm{(E)}}_{\mathrm{dis},\ell}{(\boldsymbol{\mathcal{W}})}\,$, (16)
@@
$q^{\mathrm{(E)}}_{\mathrm{dis},\ell}{(\boldsymbol{\mathcal{W}})}=\frac{1}{2}\sum_{n=\ell}^{\ell+1}\sum_{i=1}^{2}w_{n,i}S^{z}_{n,i}\,.$ (19)
With this choices for the local energy density in Eqs. (17), (18) and (19), the local energy current $j_{\ell}^{\mathrm{(E)}}$ amounts to [^2] Note that the local energy density is not uniquely defined.
$j_{\ell}^{\mathrm{(E)}}{(\boldsymbol{\mathcal{W}})}=j_{\parallel,\ell}^{\mathrm{(E)}}{(\boldsymbol{\mathcal{W}})}+j_{\perp,\ell}^{\mathrm{(E)}},$ (20)
@@
$j_{\parallel,\ell}^{\mathrm{(E)}}{(\boldsymbol{\mathcal{W}})}$ $=$ $J^{2}_{\parallel}\sum_{i=1}^{2}j_{(\ell+1,\ell-1),(i,i)}S^{z}_{\ell,i}$
@@ III.2.3 Current autocorrelation functions
With Eq. (23), the autocorrelation function of the total current operator for the disorder realization $\boldsymbol{\mathcal{W}}$, denoted by $C^{(\mu)}{(t;{\boldsymbol{\mathcal{W}}},\beta)}$, can be defined as
@@
$\overline{C^{(\mu)}}{(t;\beta)}=\mathbb{E}_{\boldsymbol{\mathcal{W}}}\quantity[C^{(\mu)}{(t;{\boldsymbol{\mathcal{W}}},\beta)}]\,.$ (25)
@@ III.2.4 Time-dependent disorder-averaged diffusion constants
Figure 4: Comparison between disorder-averaged time-dependent diffusion constants of a disordered spin-$1/2$ XX ladder as a function of time at disorder strength $W/J_{\parallel}=1.5$ and coupling ratio $J_{\perp}/J_{\parallel}=1.0$ for various numbers of rungs, $L=7,8,9,10$. The spin diffusion (purple lines) occurs at a faster rate than the energy diffusion (green lines). The opacity of each line denotes the result for a chosen $L$ — lower opacity corresponds to a smaller system, while higher opacity indicates a larger one. The arrows specify how the time-dependent diffusion constant converges with an increase in $L$.
Furthermore, as the (normalized) disorder strength $W/J_{\parallel}$ or the coupling ratio $J_{\perp}/J_{\parallel}$ increases beyond a certain threshold, the spin and energy diffusion constants tend to converge, ultimately yielding similar values.
@@ III.2.5 Disorder-averaged static susceptibilities
In the infinite-temperature and thermodynamic limits, the disorder-averaged static spin susceptibility per spin of our model, $\overline{\chi_{\mathrm{S}}}{(\beta=0)}$, turns out to be disorder-independent and becomes
@@
The disorder-averaged static energy susceptibility per spin, $\overline{\chi_{\mathrm{E}}}{(\beta=0)}$, is disorder-dependent and amounts to
@@ III.2.6 Numerical extraction of $t_{\mathrm{Th}}$
Following the computation of the disorder-averaged diffusion constant in the thermodynamic and infinite-time limits, denoted as $\overline{D^{(\mu)}}{(\beta)}$, as outlined in Eq. (27), we proceed to extract the Thouless time $t^{(\mu)}_{\mathrm{Th}}$. To achieve this, we simply substitute the obtained value of $\overline{D^{(\mu)}}{(\beta)}$ into the formula for $t^{(\mu)}_{\mathrm{Th}}$ given in Eq. (9). For all graphical and quantitative analyses of $t^{(\mu)}_{\mathrm{Th}}$ as a function of the linear-system size $L$, we use the estimated large-system-size values of the diffusion constant. Corrections must be sufficient small such that $t^{(\mu)}_{\mathrm{Th}}\propto L^{2}$ is the leading $L$-dependence.
@@
For the computation of these diffusion constants, we use the grand-canonical ensemble of the spin models, corresponding to all of the total magnetization symmetry sectors ($\expectationvalue{S^{z}}=0$), where the time evolution is carried out via the Lanczos method with dynamical quantum typicality [127, 128, 129, 130, 131, 132, 133] (see Appendix C.3 for the details). For the Lanczos algorithm, we discretize time with a time step of $\Delta t=0.01/J_{\parallel}$ and carry out a maximum of $30$ Lanczos iterations per discrete time step. To compute the disorder-averaged diffusion constants, we employ an ensemble of $M=1000$ independent disorder realizations.
We estimate $\overline{D^{(\mu)}}$ by taking an average of $\overline{D^{(\mu)}}{(t)}$ in the time interval $200\leq tJ_{\parallel}\leq 250$ for the largest two systems sizes we have ($N=18,20$). In contrast to the clean case [80], we find that taking late-time results for the estimate is more convenient because a saturation in $\overline{D^{(\mu)}}{(t)}$ emerges late in the disordered case.
@@ IV Results
IV.1 Spectral form factor and RMT time
Figure 5: Running-and-disorder-averaged spectral form factor (SFF) $\overline{K}_{\mathrm{RA}}{(t)}$ of a spin-$1/2$ XX ladder as a function of time normalized by $L^{2}$ at disorder strengths $W/J_{\parallel}=1.25$ in (a) and $W/J_{\parallel}=1.5$ in (b) and coupling ratio $J_{\perp}/J_{\parallel}=1.0$. The (real) time is obtained from the normalized time $\tau$ via $t=\frac{\tau}{2\pi}t_{\mathrm{H}}$, where $t_{\mathrm{H}}$ denotes the Heisenberg time. The color of each line corresponds to a specific number of rungs, $L$. Solid lines represent $\overline{K}_{\mathrm{RA}}{(t)}$, while black dashed lines denote the SFF of the Gaussian Orthogonal Ensemble (GOE), $K_{\mathrm{GOE}}{(t)}$. Each SFF is computed with a discrete time step in $\log_{10}$ scale, $\Delta(\log_{10}\tau)$, of $0.001/J_{\parallel}$, using a running average window in $\log_{10}$ scale, $\delta(\log_{10}\tau)$, of $0.025/J_{\parallel}$ and a Gaussian filter width factor $\eta$ of $0.5$. All SFFs are averaged over $M=5000$ disorder realizations. The random matrix theory (RMT) time $t_{\mathrm{RMT}}$ (red dots with error bars) is determined for each $L$ by varying the threshold value for the criterion function $\Delta K{(t)}$, $\epsilon_{\Delta K}$, within the range $[0.035,0.2]$.
For a disordered spin-$1/2$ XX ladder, we first present the running-and-disorder-averaged SFF, $\overline{K}_{\mathrm{RA}}{(t)}$, in Fig. 5 as a function of time normalized by the squared number of rungs, $L^{2}$, at the coupling ratio $J_{\perp}/J_{\parallel}=1.0$ and the disorder strengths $W/J_{\parallel}=1.25$ in Fig. 5(a) and $W/J_{\parallel}=1.5$ in Fig. 5(b).
Due to the choice of a large number of disorder realizations in the SFF and the exponential growth of the Hilbert-space dimension of our model with increasing the system size $N$, we compute $\overline{K}_{\mathrm{RA}}{(t)}$ for four different numbers of rungs, $L\in\{5,6,7,8\}$, where we indicate these numbers of rungs by different colors. For all system sizes we consider, we can clearly distinguish the expected three regimes in the SFF in Fig. 5. More importantly, all values of $t_{\mathrm{RMT}}$, indicated by the red dots with the error bars, align linearly along the SFF axis within the error range determined by the threshold value $\epsilon_{\Delta{K}}$ and the Gaussian filter width factor $\eta$. This implies the existence of an $L^{2}$ scaling in the RMT time, i.e., $t_{\mathrm{RMT}}{(L)}\propto L^{2}$. We notice that we could not identify $L^{2}$-scaling RMT times for small disorder strengths such as $W/J_{\parallel}\lesssim 1.0$ even for large system sizes.

### v3

@@ III.1.2 Numerical extraction of $t_{\mathrm{RMT}}$
A straightforward approach to extract $\tau_{\mathrm{RMT}}$ from $\overline{K}{(\tau)}$ involves comparing it with the SFF of a RMT eigenvalue distribution, such as the Gaussian orthogonal ensemble (GOE). This comparison leads to the following criterion function [22]:
@@
$\Delta K{(\tau_{\mathrm{RMT}})}=\epsilon_{\Delta K}\,.$ (7)
For the comparison with other relaxation timescales, the RMT time must be expressed in real units. This can be achieved by multiplying it by the real Heisenberg time $t_{\mathrm{H}}$ as follows:
@@
To eliminate the variations in the density of states $\rho{(E)}$ due to the ensemble of the various disorder realizations appearing in the SFF computations, we unfold the energy spectrum of each realization $\boldsymbol{\mathcal{W}}$, which also makes comparing the numerical results with RMT easier and thus highlights the universal RMT behavior [123, 60, 124, 125]. As a filter function, we use a Gaussian distribution with a given filter width factor $\eta$. However, for some disorder realizations, even after unfolding the energy spectrum, $\tau_{\mathrm{H}}$ may shift from $2\pi$ because of the spectral edge effects [126]. To prevent this shifting, we renormalize $\tau_{\mathrm{H}}$ by applying the filter function to the mean-level spacing of the unfolded spectrum. We explain this shift-fixing procedure in Appendix B.3 and further details regarding the unfolding and filtering procedures can be found in Appendixes B.1 and B.2, respectively.
The choices for $\epsilon_{\Delta K}$ and $\eta$ quantitatively affect the result for $t_{\mathrm{RMT}}$. However, the $L^{2}$ scaling of $t_{\mathrm{RMT}}$ observed in our model remains robust if $\epsilon_{\Delta K}$ is chosen to be sufficiently large to surpass the fluctuations that occur just prior to the onset of RMT behavior in $\overline{K}_{\mathrm{RA}}{(t)}$, and $\eta$ falls within a reasonable range, typically between $0.2$ and $1.0$. As part of our analysis, we estimate the RMT time $t_{\mathrm{RMT}}$ for the models by varying the threshold value $\epsilon_{\Delta K}$ over a range of $0.035$ to $0.2$, while keeping the parameter $\eta$ fixed at $0.5$. Generally, we search for the range of pairs $(\eta,\epsilon_{\Delta K})$ that yield the same $L$ dependence of $t_{\mathrm{RMT}}$, from which we then extract $t_{\mathrm{RMT}}$ up to this systematic uncertainty. Thus the exact prefactor of $t_{\mathrm{RMT}}$ is only fixed within these limitations.
In our results, we obtain $\overline{K}_{\mathrm{RA}}{(\tau)}$ in Eq. (6) by time-averaging $\overline{K}{(\tau)}$ with a running average window in $\log_{10}$ scale $\delta{(\log_{10}\tau)}$ of $0.025$ while the discrete time step in $\log_{10}$ scale $\Delta{(\log_{10}\tau)}$ is $0.001$.
@@ III.2 Spin- and energy-Thouless times
Our main interest is in a direct comparison of Thouless times, extracted from transport, with the RMT time. We recall that the Thouless time $t_{\mathrm{Th}}$ is the average time for a local excitation to spread through a finite system with length $L$
@@
here assuming diffusion. The symbol $\mu$ indicates the transport channel, such as spin, particle, charge, or energy, and $D^{(\mu)}$ is the diffusion constant for the corresponding transport channel in the thermodynamic limit. In Eq. (9), we divide $L$ by $2\pi$ to obtain the smallest wavenumber $k_{0}=2\pi/L$. This wavenumber also appears in the autocorrelation function of the local densities $c^{(\mu)}{(k,t)}$. For late times, this autocorrelation function decays exponentially at wavenumber $k=k_{0}$ [127], with a decay rate given by
@@ III.2.1 Transport quantities
$\begin{cases}Q^{(\mathrm{S})}=\sum_{\ell,i}S^{z}_{\ell,i}=:S^{z}\\ Q^{(\mathrm{E})}=H\end{cases}\,$, (11)
where $Q^{(\mathrm{S})}$ is the total magnetization in the z direction of the spins, and $Q^{(\mathrm{E})}$ is the total energy.
@@
$\dot{q}^{(\mu)}_{\ell}$ $=$ $i\commutator{H}{q^{(\mu)}_{\ell}}\,$, (12)
$=$ $j^{(\mu)}_{\ell-1}-j^{(\mu)}_{\ell}\,$,
where $q^{(\mu)}_{\ell}$ is the local density of transport quantity at the $\ell$ rung, so that $Q^{(\mu)}=\sum_{\ell}q_{\ell}^{(\mu)}$. From this equation, we can derive expressions for the local current operator $j^{(\mu)}_{\ell}$, which are essential for characterizing the transport properties of the system. As expected, these current operators may differ depending on the choice of local density.
III.2.2 Local densities of transport quantities and local current operators
@@
[…]
Similarly, the local energy current $j_{\ell}^{\mathrm{(E)}}$ can be obtained by selecting the total energy as the transport quantity, i.e., $Q^{\mathrm{(E)}}=\sum_{\ell}q^{\mathrm{(E)}}_{\ell}$, where $q^{\mathrm{(E)}}_{\ell}$ is the local energy density. For our model, we express $q^{\mathrm{(E)}}_{\ell}$ as the sum of three terms
$q^{\mathrm{(E)}}_{\ell}\coloneqq q^{\mathrm{(E)}}_{\parallel,\ell}+q^{\mathrm{(E)}}_{{}_{\perp},\ell}+q^{\mathrm{(E)}}_{\mathrm{dis},\ell}\,$, (15)
@@
$q^{\mathrm{(E)}}_{\mathrm{dis},\ell}=\frac{1}{2}\sum_{n=\ell}^{\ell+1}\sum_{i=1}^{2}w_{n,i}S^{z}_{n,i}\,.$ (18)
With these choices for the local energy density in Eqs. (16)–(18), the local energy current $j_{\ell}^{\mathrm{(E)}}$ amounts to [^2] Note that the local energy density is not uniquely defined.
$j_{\ell}^{\mathrm{(E)}}=j_{\parallel,\ell}^{\mathrm{(E)}}+j_{\perp,\ell}^{\mathrm{(E)}},$ (19)
@@
$j_{\parallel,\ell}^{\mathrm{(E)}}$ $=$ $J^{2}_{\parallel}\sum_{i=1}^{2}j_{(\ell+1,\ell-1),(i,i)}S^{z}_{\ell,i}$
@@ III.2.3 Current autocorrelation functions
With Eq. (22), the autocorrelation function of the total current operator for the disorder realization $\boldsymbol{\mathcal{W}}$, denoted by $C^{(\mu)}{(t;\beta)}$, can be defined as
@@
$\overline{C^{(\mu)}}{(t;\beta)}=\mathbb{E}_{\boldsymbol{\mathcal{W}}}\quantity[C^{(\mu)}{(t;\beta)}]\,.$ (24)
@@ III.2.4 Time-dependent disorder-averaged diffusion constants
Figure 4: Comparison between disorder-averaged time-dependent diffusion constants of a disordered spin-$1/2$ XX ladder as a function of time at disorder strength $W/J_{\parallel}=1.5$ and coupling ratio $J_{\perp}/J_{\parallel}=1.0$ for various numbers of rungs $L=7,8,9,10$. The spin diffusion (purple lines) occurs at a faster rate than the energy diffusion (green lines). The opacity of each line denotes the result for a chosen LL—lower opacity corresponds to a smaller system, while higher opacity indicates a larger one. The arrows specify how the time-dependent diffusion constant converges with an increase in $L$.
Furthermore, as the (normalized) disorder strength $W/J_{\parallel}$ or the coupling ratio $J_{\perp}/J_{\parallel}$ increases beyond a certain threshold, the spin- and energy-diffusion constants tend to converge to the same result that is independent of $J_{\perp}$.
@@ III.2.5 Disorder-averaged static susceptibilities
In the infinite-temperature and thermodynamic limits, the disorder-averaged static spin susceptibility per spin of our model $\overline{\chi_{\mathrm{S}}}{(\beta=0)}$ turns out to be disorder independent and becomes
@@
The disorder-averaged static energy susceptibility per spin $\overline{\chi_{\mathrm{E}}}{(\beta=0)}$ is disorder dependent and amounts to
@@ III.2.6 Numerical extraction of $t_{\mathrm{Th}}$
Following the computation of the disorder-averaged diffusion constant in the thermodynamic and infinite-time limits, denoted as $\overline{D^{(\mu)}}{(\beta)}$, as outlined in Eq. (26), we proceed to extract the Thouless time $t^{(\mu)}_{\mathrm{Th}}$. To achieve this, we simply substitute the obtained value of $\overline{D^{(\mu)}}{(\beta)}$ into the formula for $t^{(\mu)}_{\mathrm{Th}}$ given in Eq. (9). For all graphical and quantitative analyses of $t^{(\mu)}_{\mathrm{Th}}$ as a function of the linear system size $L$, we use the estimated large-system-size values of the diffusion constant. Corrections must be sufficient small such that $t^{(\mu)}_{\mathrm{Th}}\propto L^{2}$ is the leading $L$ dependence.
@@
For the computation of these diffusion constants, we use the grand-canonical ensemble of the spin models, corresponding to all of the total magnetization symmetry sectors ($\expectationvalue{S^{z}}=0$), where the time evolution is carried out via the Lanczos method with dynamical quantum typicality (DQT) [129, 130, 131, 132, 133, 134, 135] (see Appendix C.3 for the details). For the Lanczos algorithm, we discretize time with a time step of $\Delta t=0.01/J_{\parallel}$ and carry out a maximum of $30$ Lanczos iterations. To compute the disorder-averaged diffusion constants, we employ an ensemble of $M=1000$ independent disorder realizations.
We estimate $\overline{D^{(\mu)}}$ by taking an average of $\overline{D^{(\mu)}}{(t)}$ in the time interval $200\leq tJ_{\parallel}\leq 250$ for the largest two systems sizes we have ($N=18,20$). In contrast with the clean case [79], we find that taking late-time results for the estimate is more convenient because a saturation in $\overline{D^{(\mu)}}{(t)}$ emerges late in the disordered case.
@@ IV Results
IV.1 SFF and RMT time
Figure 5: Running-and-disorder-averaged SFF $\overline{K}_{\mathrm{RA}}{(t)}$ of a spin-$1/2$ XX ladder as a function of time normalized by $L^{2}$ at disorder strengths (a) $W/J_{\parallel}=1.25$ and (b) $W/J_{\parallel}=1.5$ and coupling ratio $J_{\perp}/J_{\parallel}=1.0$. The (real) time is obtained from the normalized time $\tau$ via $t=\frac{\tau}{2\pi}t_{\mathrm{H}}$, where $t_{\mathrm{H}}$ denotes the Heisenberg time. The color of each line corresponds to a specific number of rungs, $L$. Solid lines represent $\overline{K}_{\mathrm{RA}}{(t)}$, while black dashed lines denote the SFF of the GOE $K_{\mathrm{GOE}}{(t)}$. Each SFF is computed with a discrete time step in $\log_{10}$ scale $\Delta(\log_{10}\tau)$ of $0.001$ using a running average window in $\log_{10}$ scale $\delta(\log_{10}\tau)$ of $0.025$ and a Gaussian filter width factor $\eta$ of $0.5$. All SFFs are averaged over $M=5000$ disorder realizations. The RMT time $t_{\mathrm{RMT}}$ (red dots with error bars) is determined for each $L$ by varying the threshold value for the criterion function $\Delta K{(t)}$, $\epsilon_{\Delta K}$, within the range $[0.035,0.2]$.
For a disordered spin-$1/2$ XX ladder, we first present the running-and-disorder-averaged SFF $\overline{K}_{\mathrm{RA}}{(t)}$ in Fig. 5 as a function of time normalized by the squared number of rungs $L^{2}$ at the coupling ratio $J_{\perp}/J_{\parallel}=1.0$ and the disorder strengths $W/J_{\parallel}=1.25$ in Fig. 5(a) and $W/J_{\parallel}=1.5$ in Fig. 5(b).
Due to the choice of a large number of disorder realizations in the SFF and the exponential growth of the Hilbert-space dimension of our model with increasing the system size $N$, we compute $\overline{K}_{\mathrm{RA}}{(t)}$ for four different numbers of rungs $L\in\{5,6,7,8\}$, where we indicate these numbers of rungs by different colors. For all system sizes we consider, we can clearly distinguish the expected three regimes in the SFF in Fig. 5. More importantly, all values of $t_{\mathrm{RMT}}$, indicated by the red dots with the error bars, align linearly along the SFF axis within the error range determined by the threshold value $\epsilon_{\Delta{K}}$ and the Gaussian filter width factor $\eta$. This implies the existence of an $L^{2}$ scaling in the RMT time, i.e., $t_{\mathrm{RMT}}{(L)}\propto L^{2}$. We notice that we could not identify $L^{2}$-scaling of RMT times for small disorder strengths such as $W/J_{\parallel}\lesssim 1.0$ even for large system sizes.